# Rotatable and Movable Antenna-Enabled Near-Field Integrated Sensing and Communication

### v3

@@ Abstract
Index Terms:
Integrated sensing and communications (ISAC), near-field, particle swarm optimization (PSO), rotatable and movable antenna (RMA).
@@ I Introduction
Integrated Sensing and Communication (ISAC) is a promising technology in the evolution of the sixth generation (6G) wireless systems, which converges the traditionally separate domains of communication and sensing into a unified framework [1]. Due to its ability to efficiently allocate time, frequency, power, and hardware resources for simultaneous communication and sensing tasks [2], ISAC has been officially approved by IMT-2030 as one of six decisive usage scenarios for 6G networks and attracted increasing research attention. To further enhance the performance of sensing and communication, ISAC has been combined with multiple-antenna technologies, such as multiple-input multiple-output (MIMO) [3], to exploit additional spatial degrees of freedom (DoFs) and achieve greater array gains for both communication and sensing tasks. These improvements contribute to higher network throughput and improved sensing resolution. In general, the performance gains offered by antenna arrays scale approximately linearly with the aperture size. This insight has inspired the development of large-aperture array architectures, including reconfigurable intelligent surfaces (RIS) [4, 5], dynamic metasurfaces [6], holographic MIMO [7], and other related technologies, which have been widely adopted in ISAC systems.
@@ I-A Prior Works
Benefiting from near-field propagation, near-field ISAC can obtain accurate resolution in spatial sensing, such as joint sensing of targets’ angle and distance information, while ensuring high communication performance. Nevertheless, most near-field ISAC systems are based on conventional fixed-position antennas (FPAs). Such deployment limits the diversity and spatial multiplexing performance of MIMO systems because the channel variation in the spatial field is not fully utilized [11].
To overcome the inherent limitations of FPAs, movable antennas (MAs) and six-dimensional MAs (6DMAs) are introduced to wireless networks, which can fully exploit the spatial DoFs of antenna arrays in communication or sensing scenarios [12, 13, 14, 15]. MA-enabled systems can dynamically adjust the positions of antennas, thereby reshaping the channel conditions to boost communication performance, or reconfiguring the geometric properties of the antenna arrays to enhance sensing capability [16]. Building upon MAs, 6DMA leverages the DoFs in both the three-dimensional (3D) positioning and rotation of antenna planar. Due to the advantageous properties of MA and 6DMA, there has recently been extensive research on various wireless sensing and ISAC systems enabled by these technologies [16, 17, 18, 19, 20, 21]. The authors in [16] investigated a MA-enabled ISAC system with RIS and maximized the minimum beampattern gain by jointly optimizing the transmit beamforming, the positions of MAs, and the phase shifts of RIS. In addition, the authors in [17] studied the Cramér-Rao bound (CRB) minimization problem for a MA-enabled ISAC system. Based on existing research on the MA-enabled far-field ISAC systems [18, 19], the authors in [20] investigated the application of MAs in a dual-functional full-duplex ISAC system with near-field channel model. Moreover, the authors in [21] proposed a wireless sensing system enabled by 6DMA and minimized the CRB for estimating the directions of arrival by jointly optimizing the 6DMAs’ positions and rotations at the base station (BS), comparing it with MAs for directive and isotropic antenna radiation patterns.
@@ I-C Organization and Notations
Notations: Throughout this paper, scalars, vectors, and matrices are denoted by non-bold, bold lower-case, and bold upper-case letters, respectively. For the vector $\mathbf{a}$, $[\mathbf{a}]_{i}$, $\mathbf{a}^{\mathsf{T}}$, $\mathbf{a}^{*}$, and $\mathbf{a}^{\mathsf{H}}$ denote the $i$-th entry, transpose, conjugate, and conjugate transpose of $\mathbf{a}$, respectively. The notations $|\mathbf{a}|$ and $\|\mathbf{a}\|$ denote the magnitude and norm of scalar $\mathbf{a}$ and vector $\mathbf{a}$, respectively. $\mathbf{A}\succeq\mathbf{0}$ means that matrix $\mathbf{A}$ is positive semidefinite; $\mathbf{A}\succeq\mathbf{B}$ means that $\mathbf{A-B}\succeq\mathbf{0}$; $\text{rank}(\mathbf{A})$ and $\text{tr}(\mathbf{A})$ denote the rank and trace of matrix $\mathbf{A}$, respectively; ${}\Re\{\cdot\}$ and ${}\Im\{\cdot\}$ denote the real and imaginary component of a complex number, respectively. The $N\times N$ identity matrix is denoted by $\mathbf{I}_{N}$. The set $\mathbb{C}^{N\times M}$ and $\mathbb{R}^{N\times M}$ stands for the space of ${N\times M}$ complex and real matrices, respectively. Finally, $\mathcal{CN}(\mu,\mathbf{X})$ is used to denote the circularly symmetric complex Gaussian distribution with mean $\mu$ and covariance matrix $\mathbf{X}$.
@@ II-A RMA-BS Model
$\mathbf{q}_{n_{\text{t}}}^{\text{t}}$ $=[0,y_{n_{\text{t}}}^{\text{t}},z_{n_{\text{t}}}^{\text{t}}]^{\mathsf{T}}\in\mathcal{C}^{\text{t}},$ (1)
$\mathbf{q}_{n_{\text{r}}}^{\text{r}}$ $=[0,y_{n_{\text{r}}}^{\text{r}},z_{n_{\text{r}}}^{\text{r}}]^{\mathsf{T}}\in\mathcal{C}^{\text{r}},$ (2)
where $\mathcal{C}^{\text{t}}$ and $\mathcal{C}^{\text{r}}$ denote the square moving regions of size $D\times D$ with the center points at the origin of corresponding local coordinates.
@@
$\bm{\varpi}^{\text{t}}$ $=[\alpha^{\text{t}},\beta^{\text{t}},\gamma^{\text{t}}]^{\mathsf{T}},$ (3)
$\bm{\varpi}^{\text{r}}$ $=[\alpha^{\text{r}},\beta^{\text{r}},\gamma^{\text{r}}]^{\mathsf{T}},$ (4)
[…]
@@ II-B Channel Model
[…]
@@
[…]
@@ II-E Communication Performance Metric: Communication Rate
In this paper, we assume that the perfect channel state information (CSI) of communication channels is perfectly known at the BS via a proper channel estimation mechanism [29]. From (24), the received SINR at the $k$-th UT can be expressed as follows:
@@ III-C Optimization of RMAs’ Positions and Rotations
where ${y^{\text{p}}}^{(0)}_{m,n_{\text{p}}},{z^{\text{p}}}^{(0)}_{m,n_{\text{p}}}\sim\mathcal{U}(-\frac{D}{2},\frac{D}{2})$; ${\alpha^{\text{p}}}^{(0)}_{m},{\beta^{\text{p}}}^{(0)}_{m},{\gamma^{\text{p}}}^{(0)}_{m}\sim\mathcal{U}(0,2\pi)$ for $m\in[1,M]$, $n_{\text{p}}\in[1,N_{\text{p}}]$. Then we randomly initialize each particle’s velocity vector $\mathbf{v}^{(0)}_{m}\in\mathbb{R}^{(2N_{\text{r}}+2N_{\text{t}}+6)\times 1}$ and define $\bm{\Xi}^{(0)}=\{\bm{\xi}^{(0)}_{1},\dots,\bm{\xi}^{(0)}_{M}\}$, $\mathbf{V}^{(0)}=\{\mathbf{v}^{(0)}_{1},\dots,\mathbf{v}^{(0)}_{M}\}$. Subsequently, we set the initial local best position of M particles as $\bm{\xi}_{m,{\text{l}best}}=\bm{\xi}_{m}^{(0)}$ and select the particle with the best fitness as the global particle $\bm{\xi}_{\text{g}best}$ of the swarm. The fitness of the particles is defined in the subsequent text.
@@
1:Initialize $i=0$,$\mathbf{W},\mathbf{R}_{0},\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$.
@@
5:     Calculate $f_{0}={\text{tr}}\left({\text{CRB}}\left(\mathbf{R}_{x},\mathbf{G},\sigma_{0}^{2}\right)\right)$ by (27).
7:     Set the local best positions $\bm{\xi}_{m,{\text{l}best}}=\bm{\xi}_{m}^{(0)}$ for $m\in[1,M]$ and the global best positions $\bm{\xi}_{{\text{g}best}}=\operatorname*{argmin}_{\bm{\xi}_{m}^{(0)}}\{\mathcal{F}(\bm{\xi}_{1}^{(0)}),\dots,\mathcal{F}(\bm{\xi}_{M}^{(0)})\}$
8:     for $\tau=1$ to $\tau_{\max}$ do
9:         Update the inertia parameter $\omega$ according to (41).
10:         for $m=1$ to M do
11:              Update the velocity and position of the $m$-th particle according to (39) and (40), respectively.
12:              Calculate the fitness value of the $m$-th particle, i.e., $\mathcal{F}(\bm{\xi}_{m}^{(\tau)})$, according to (44).
13:              if $\mathcal{F}(\bm{\xi}_{m}^{(\tau)})<\mathcal{F}(\bm{\xi}_{m,\text{lbest}})$ then
14:                  Update $\bm{\xi}_{m,\text{lbest}}=\bm{\xi}_{m}^{(\tau)}$.
15:              end if
16:              if $\mathcal{F}(\bm{\xi}_{m}^{(\tau)})<\mathcal{F}(\bm{\xi}_{\text{gbest}})$ then
17:                  Update $\bm{\xi}_{\text{gbest}}=\bm{\xi}_{m}^{(\tau)}$.
18:              end if
@@
20:     end for
21:     if $\mathcal{F}(\bm{\xi}_{\text{gbest}})<f_{0}$ then
23:     end if
24:     Set $i=i+1$.
25:until Convergence or $i>I_{\max}$
@@ III-D Convergence and Complexity Analysis
Then we analyze the computational complexity of the proposed algorithm. For each iteration, the complexity for updating $\mathbf{W}$ and $\mathbf{R}_{0}$ via the interior method is in order of $\mathcal{O}(K^{6.5}M^{6.5}{\text{log}}\epsilon^{-1})$ with the given solution accuracy $\epsilon$ [33]. And the complexity for updating $\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$ via the PSO method is in order of $\mathcal{O}\left(\tau_{\max}M(2N_{\text{t}}+2N_{\text{r}}+6+\text{log}M)\right)$. Therefore, the total computational complexity of Algorithm 1 is $\mathcal{O}\left(I_{\text{AO}}(K^{6.5}M^{6.5}{\text{log}}\epsilon^{-1}+\tau_{\max}M(2N_{\text{t}}+2N_{\text{r}}+6+\text{log}M))\right)$ , with $I_{\text{AO}}$ denoting the resultant iteration number of the AO-based algorithm.
@@ IV-D Optimization of RMAs’ Positions and Rotations
1:Initialize $i=0$,$\mathbf{W},\mathbf{R}_{0},\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$.
@@
6:     Calculate $e_{0}=f_{2}(\mathbf{W},\mathbf{R}_{0},\mathbf{H},\bm{\rho},\bm{\nu})$.
7:     Initialize the M particles with positions $\bm{\Xi}^{(0)}$ and velocity $\mathbf{V}^{(0)}$.
8:     Update $\mu_{4}$ by (60).
9:     Set the local best positions $\bm{\xi}_{m,{\text{l}best}}=\bm{\xi}_{m}^{(0)}$ for $m\in[1,M]$ and the global best positions $\bm{\xi}_{{\text{g}best}}=\operatorname*{argmax}_{\bm{\xi}_{m}^{(0)}}\{\mathcal{E}(\bm{\xi}_{1}^{(0)}),\dots,\mathcal{E}(\bm{\xi}_{M}^{(0)})\}$
10:     for $\tau=1$ to $\tau_{\max}$ do
11:         Update the inertia parameter $\omega$ according to (41).
12:         for $m=1$ to M do
13:              Update the velocity and position of the $m$-th particle according to (39) and (40), respectively.
14:              Calculate the fitness value of the $m$-th particle, i.e., $\mathcal{E}(\bm{\xi}_{m}^{(\tau)})$, according to (58).
15:              if $\mathcal{E}(\bm{\xi}_{m}^{(\tau)})>\mathcal{E}(\bm{\xi}_{m,\text{lbest}})$ then
16:                  Update $\bm{\xi}_{m,\text{lbest}}=\bm{\xi}_{m}^{(\tau)}$.
17:              end if
18:              if $\mathcal{E}(\bm{\xi}_{m}^{(\tau)})>\mathcal{E}(\bm{\xi}_{\text{gbest}})$ then
19:                  Update $\bm{\xi}_{\text{gbest}}=\bm{\xi}_{m}^{(\tau)}$.
20:              end if
@@
22:     end for
23:     if $\mathcal{E}(\bm{\xi}_{\text{gbest}})>e_{0}$ then
24:         Update $\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$ according to $\bm{\xi}_{\text{gbest}}$.
25:     end if
26:     Set $i=i+1$.
27:until Convergence or $i>I_{\max}$

### v4

@@ I Introduction
Integrated Sensing and Communication (ISAC) is a promising technology in the evolution of the sixth generation (6G) wireless systems, which converges the traditionally separate domains of communication and sensing into a unified framework [1]. Due to its ability to efficiently allocate time, frequency, power, and hardware resources for simultaneous communication and sensing tasks [2], ISAC has been officially approved by IMT-2030 as one of six decisive usage scenarios for 6G networks and attracted increasing research attention. To further enhance the performance of sensing and communication, ISAC has been combined with multiple-antenna technologies, such as multiple-input multiple-output (MIMO) [3, 4], to exploit additional spatial degrees of freedom (DoFs) and achieve greater array gains for both communication and sensing tasks. These improvements contribute to higher network throughput and improved sensing resolution. In general, the performance gains offered by antenna arrays scale approximately linearly with the aperture size. This insight has inspired the development of large-aperture array architectures, including reconfigurable intelligent surfaces (RIS) [5, 6], dynamic metasurfaces [7], holographic MIMO [8], and other related technologies, which have been widely adopted in ISAC systems.
@@ I-A Prior Works
Benefiting from near-field propagation, near-field ISAC can obtain accurate resolution in spatial sensing, such as joint sensing of targets’ angle and distance information, while ensuring high communication performance. Recent studies have demonstrated the feasibility and of acquiring three-dimensional (3D) information through near-field sensing and ISAC techniques [12, 13]. Nevertheless, most near-field ISAC systems are based on conventional fixed-position antennas (FPAs). Such deployment limits the diversity and spatial multiplexing performance of MIMO systems because the channel variation in the spatial field is not fully utilized [14].
To overcome the inherent limitations of FPAs, movable antennas (MAs) and six-dimensional MAs (6DMAs) are introduced to wireless networks, which can fully exploit the spatial DoFs of antenna arrays in communication or sensing scenarios [15, 16, 17, 18]. MA-enabled systems can dynamically adjust the positions of antennas, thereby reshaping the channel conditions to boost communication performance, or reconfiguring the geometric properties of the antenna arrays to enhance sensing capability [19]. Building upon MAs, 6DMA leverages the DoFs in both the 3D positioning and rotation of antenna planar. Due to the advantageous properties of MA and 6DMA, there has recently been extensive research on various wireless sensing and ISAC systems enabled by these technologies [19, 20, 21, 22, 23, 24]. The authors in [19] investigated a MA-enabled ISAC system with RIS and maximized the minimum beampattern gain by jointly optimizing the transmit beamforming, the positions of MAs, and the phase shifts of RIS. In addition, the authors in [20] studied the Cramér-Rao bound (CRB) minimization problem for a MA-enabled ISAC system. Based on existing research on the MA-enabled far-field ISAC systems [21, 22], the authors in [23] investigated the application of MAs in a dual-functional full-duplex ISAC system with near-field channel model. Moreover, the authors in [24] proposed a wireless sensing system enabled by 6DMA and minimized the CRB for estimating the directions of arrival by jointly optimizing the 6DMAs’ positions and rotations at the base station (BS), comparing it with MAs for directive and isotropic antenna radiation patterns.
@@ I-C Organization and Notations
Notations: Throughout this paper, scalars, vectors, and matrices are denoted by non-bold, bold lower-case, and bold upper-case letters, respectively. For the vector $\mathbf{a}$, $\mathbf{a}]_{i}$, $\mathbf{a}^{\mathsf{T}}$, $\mathbf{a}^{*}$, and $\mathbf{a}^{\mathsf{H}}$ denote the $i$-th entry, transpose, conjugate, and conjugate transpose of $\mathbf{a}$, respectively. The notations |a||{a}| and $\|\mathbf{b}\|$ denote the magnitude and norm of scalar $a$ and vector $\mathbf{b}$, respectively. $\mathbf{A}\succeq\mathbf{0}$ means that matrix $\mathbf{A}$ is positive semidefinite; $\mathbf{A}\succeq\mathbf{B}$ means that $\mathbf{A-B}\succeq\mathbf{0}$; $\text{rank}(\mathbf{A})$ and $\text{tr}(\mathbf{A})$ denote the rank and trace of matrix $\mathbf{A}$, respectively; $\Re\{\cdot\}$ and $\Im\{\cdot\}$ denote the real and imaginary component of a complex number, respectively. The $N\times N$ identity matrix is denoted by $\mathbf{I}_{N}$. The set $\mathbb{C}^{N\times M}$ and $\mathbb{R}^{N\times M}$ stands for the space of $N\times M$} complex and real matrices, respectively. Finally, $\mathcal{CN}(\mu,\mathbf{X})$ is used to denote the circularly symmetric complex Gaussian distribution with mean $\mu$ and covariance matrix $\mathbf{X}$.
@@ II-A RMA-BS Model
$\mathbf{q}_{n_{\text{p}}}^{\text{p}}$ $=[0,y_{n_{\text{p}}}^{\text{p}},z_{n_{\text{p}}}^{\text{p}}]^{\mathsf{T}}\in\mathcal{C}^{\text{p}},$ (1)
where ${\text{p}}\in\{\text{t,r}\}$; $\mathcal{C}^{\text{t}}$ and $\mathcal{C}^{\text{r}}$ denote the square moving regions of size $D\times D$ with the center points at the origin of corresponding local coordinates.
@@
$\bm{\varpi}^{\text{p}}$ $=[\alpha^{\text{p}},\beta^{\text{p}},\gamma^{\text{p}}]^{\mathsf{T}},$ (2)
where $\alpha^{\text{p}}\in[0,2\pi)$, $\beta^{\text{p}}\in[0,2\pi)$, and $\gamma^{\text{p}}\in[0,2\pi)$ denote the rotation angles with respect to (w.r.t.) the $x$-axis, $y$-axis, and $z$-axis, respectively, with ${\text{p}}\in\{\text{t,r}\}$.
@@ II-B Channel Model
$|{h}^{\text{p}}_{n_{\text{p}},k}|^{2}=\int_{\mathcal{S}_{n_{\text{p}}}}\mathcal{L}^{\text{p}}(\mathbf{p}_{k},\mathbf{s})\mathcal{G}^{\text{p}}(\mathbf{p}_{k},\mathbf{s})d\mathbf{s},$ (13)
@@
for ${\text{p}}={\text{t}}$ and r, respectively. As a result, the near-field channel between the $k$-th UT/target and the $n_{\text{p}}$ antenna can be derived as
@@ II-E Communication Performance Metric: Communication Rate
In this paper, we assume that the perfect channel state information (CSI) of communication channels is perfectly known at the BS via a proper channel estimation mechanism [32, 33]. From (22), the received SINR at the $k$-th UT can be expressed as follows:
@@ III-C Optimization of RMAs’ Positions and Rotations
where ${y^{\text{p}}}^{(0)}_{1,n_{\text{p}}}=y^{\text{p}}_{n_{\text{p}}},{z^{\text{p}}}^{(0)}_{1,n_{\text{p}}}=z^{\text{p}}_{n_{\text{p}}}$, ${\alpha^{\text{p}}}^{(0)}_{1}=\alpha^{\text{p}},{\beta^{\text{p}}}^{(0)}_{1}=\beta^{\text{p}},{\gamma^{\text{p}}}^{(0)}_{1}=\gamma^{\text{p}}$, for $n_{\text{p}}\in[1,N_{\text{p}}]$; ${y^{\text{p}}}^{(0)}_{m,n_{\text{p}}},{z^{\text{p}}}^{(0)}_{m,n_{\text{p}}}\sim\mathcal{U}(-\frac{D}{2},\frac{D}{2})$, ${\alpha^{\text{p}}}^{(0)}_{m},{\beta^{\text{p}}}^{(0)}_{m},{\gamma^{\text{p}}}^{(0)}_{m}\sim\mathcal{U}(0,2\pi)$ for $m\in[2,M]$, $n_{\text{p}}\in[1,N_{\text{p}}]$, with ${\text{p}}\in\{\text{t,r}\}$. Then we randomly initialize each particle’s velocity vector $\mathbf{v}^{(0)}_{m}\in\mathbb{R}^{(2N_{\text{r}}+2N_{\text{t}}+6)\times 1}$ and define $\bm{\Xi}^{(0)}=\{\bm{\xi}^{(0)}_{1},\dots,\bm{\xi}^{(0)}_{M}\}$, $\mathbf{V}^{(0)}=\{\mathbf{v}^{(0)}_{1},\dots,\mathbf{v}^{(0)}_{M}\}$. Subsequently, we set the initial local best position of M particles as $\bm{\xi}_{m,{\text{lbest}}}=\bm{\xi}_{m}^{(0)}$ and select the particle with the best fitness as the global particle $\bm{\xi}_{\text{gbest}}$ of the swarm. The fitness of the particles is defined in the subsequent text.
@@
[…]
@@
5:  Initialize the M particles with positions $\bm{\Xi}^{(0)}$ and velocity $\mathbf{V}^{(0)}$.
6:  Set the local best positions $\bm{\xi}_{m,{\text{lbest}}}=\bm{\xi}_{m}^{(0)}$ for $m\in[1,M]$ and the global best positions $\bm{\xi}_{{\text{gbest}}}=\operatorname*{argmin}_{\bm{\xi}_{m}^{(0)}}\{\mathcal{F}(\bm{\xi}_{1}^{(0)}),\dots,\mathcal{F}(\bm{\xi}_{M}^{(0)})\}$
7:  for $\tau=1$ to $\tau_{\max}$ do
8:   Update the inertia parameter $\omega$ according to (39).
9:   for $m=1$ to M do
10:     Update the velocity and position of the $m$-th particle according to (37) and (38), respectively.
11:     Calculate the fitness value of the $m$-th particle, i.e., $\mathcal{F}(\bm{\xi}_{m}^{(\tau)})$, according to (42).
12:     if $\mathcal{F}(\bm{\xi}_{m}^{(\tau)})<\mathcal{F}(\bm{\xi}_{m,\text{lbest}})$ then
13:      Update $\bm{\xi}_{m,\text{lbest}}=\bm{\xi}_{m}^{(\tau)}$.
14:     end if
15:     if $\mathcal{F}(\bm{\xi}_{m}^{(\tau)})<\mathcal{F}(\bm{\xi}_{\text{gbest}})$ then
16:      Update $\bm{\xi}_{\text{gbest}}=\bm{\xi}_{m}^{(\tau)}$.
17:     end if
18:   end for
@@
20:  Update $\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$ according to $\bm{\xi}_{\text{gbest}}$.
21:  Set $i=i+1$.
22:until Convergence or $i>I_{\max}$
@@ III-D Convergence and Complexity Analysis
Then we analyze the computational complexity of the proposed algorithm. For each iteration, the complexity for updating $\mathbf{W}$ and $\mathbf{R}_{0}$ via the interior method is in order of $\mathcal{O}(K^{6.5}N_{\text{t}}^{6.5}{\text{log}}\epsilon^{-1})$ with the given solution accuracy $\epsilon$ [28]. And the complexity for updating $\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$ via the PSO method is in order of $\mathcal{O}\left(\tau_{\max}M(2N_{\text{t}}+2N_{\text{r}}+6+\text{log}M)\right)$. Therefore, the total computational complexity of Algorithm 1 is $\mathcal{O}\left(I_{\text{AO}}(K^{6.5}N^{6.5}_{\text{t}}{\text{log}}\epsilon^{-1}+\tau_{\max}M(2N_{\text{t}}+2N_{\text{r}}+6+\text{log}M))\right)$ , with $I_{\text{AO}}$ denoting the resultant iteration number of the AO-based algorithm.
@@ IV-C Optimization of Transmit Beamforming
Note that (46b) has a necessary and sufficient condition in the form of (30b), which does not guarantee a sufficiently small CRB for each element of the target’s position information. In this paper, we adopt the stricter condition (52) to ensure the sensing accuracy and reliability.
@@ IV-D Optimization of RMAs’ Positions and Rotations
[…]
@@
7:  Update $\mu_{4}$ by (58).
8:  Set the local best positions $\bm{\xi}_{m,{\text{lbest}}}=\bm{\xi}_{m}^{(0)}$ for $m\in[1,M]$ and the global best positions $\bm{\xi}_{{\text{gbest}}}=\operatorname*{argmax}_{\bm{\xi}_{m}^{(0)}}\{\mathcal{E}(\bm{\xi}_{1}^{(0)}),\dots,\mathcal{E}(\bm{\xi}_{M}^{(0)})\}$
9:  for $\tau=1$ to $\tau_{\max}$ do
10:   Update the inertia parameter $\omega$ according to (39).
11:   for $m=1$ to M do
12:     Update the velocity and position of the $m$-th particle according to (37) and (38), respectively.
13:     Calculate the fitness value of the $m$-th particle, i.e., $\mathcal{E}(\bm{\xi}_{m}^{(\tau)})$, according to (56).
14:     if $\mathcal{E}(\bm{\xi}_{m}^{(\tau)})>\mathcal{E}(\bm{\xi}_{m,\text{lbest}})$ then
15:      Update $\bm{\xi}_{m,\text{lbest}}=\bm{\xi}_{m}^{(\tau)}$.
16:     end if
17:     if $\mathcal{E}(\bm{\xi}_{m}^{(\tau)})>\mathcal{E}(\bm{\xi}_{\text{gbest}})$ then
18:      Update $\bm{\xi}_{\text{gbest}}=\bm{\xi}_{m}^{(\tau)}$.
19:     end if
20:   end for
@@
23:  Set $i=i+1$.
24:until Convergence or $i>I_{\max}$